# Mind the Confidence Gap: Overconfidence, Calibration, and Distractor Effects in Large Language Models

### v1

@@ 2 Experimental Setup
To evaluate performance, we measure correctly answered questions for both variations ($\mathcal{N}$ and $\mathcal{D}$). For calibration assessment, we use Expected Calibration Error (ECE) to quantify the misalignment between a model’s predicted confidence and actual accuracy. A well-calibrated model produces confidence estimates that closely match its true correctness, with an ECE of zero indicating perfect calibration. Following Naeini et al. (2015), we compute ECE using empirical binning (bin size 0.1) to ensure a robust measurement of miscalibration. Additionally, we introduce two metrics, $\mathcal{D}_{helped}$ and $\mathcal{D}_{harmed}$, which measure the number of instances where adding options either improved or worsened the model’s performance.
3 Results Analysis and Discussion
Performance across LLMs
Results indicate a clear performance gap between bigger and smaller LLMs, with bigger models (e.g., GPT-4o) significantly outperforming smaller ones in open-ended question answering (Table 1). However, when provided with structured answer choices, smaller models exhibit disproportionately higher gains, while bigger models experience relative increase in distractor-induced errors ($\mathcal{D}_{harmed}$). This suggests that bigger models rely more heavily on associative recall and self-generated confidence heuristics, which can be misleading when plausible but incorrect distractors are introduced.
For instance, GPT-4o’s accuracy jumps from 35.14% ($\mathcal{N}$) to 73.42% ($\mathcal{D}$), while LLaMA3-8B exhibits an even more dramatic improvement from 4.79% to 44.01%. The fact that smaller models show a greater relative benefit from answer choices suggests that their decision-making relies more on structured elimination strategies rather than independent fact retrieval. This finding highlights a key consideration for practical deployment: while bigger models can be more self-reliant, their confidence in unaided settings does not necessarily translate into robustness when faced with misleading choices.
Calibration and Confidence Patterns
A well-calibrated model should align its confidence scores with actual accuracy—meaning that if a model expresses 80% confidence in an answer, it should be correct roughly 80% of the time. Our study reveals that bigger models, while better calibrated overall, still exhibit systematic overconfidence in open-ended tasks. The ECE for GPT-4o decreases significantly in the $\mathcal{D}$ setting, demonstrating that answer choices improve confidence alignment by providing external constraints. However, smaller models, while benefiting from options, still fail to estimate their uncertainty meaningfully, as evidenced by consistently high ECE values.
Figure 2 shows that calibration behavior varies significantly across models. While GPT-4o is relatively well-calibrated at higher confidence levels (70–100%), minor overconfidence remains. In contrast, smaller models frequently overestimate their accuracy, assigning high confidence to incorrect answers, particularly in the 80–100% range. Additionally, all LLMs at lower confidence levels (20–40%) show inconsistencies in uncertainty estimation, sometimes underestimating correctness but not necessarily engaging in random guessing. This distinction has significant implications in real-world applications: in AI-driven tutoring or medical diagnosis, overconfidence in incorrect answers can lead to misleading recommendations. Calibration-aware interventions—such as confidence regularization or explicit uncertainty weighting—are necessary to mitigate these risks.
Performance Across Question Types
To analyze calibration weaknesses, we evaluate performance across different question types (Figure 3). Person-based queries are most challenging, likely due to name ambiguities and inherent variability in names, overlapping roles, and contextual dependencies that require deeper reasoning beyond surface-level pattern matching. LLMs frequently confuse historical figures with similar names, but providing structured answer choices significantly improves accuracy, suggesting that explicit disambiguation helps mitigate uncertainty in this category. In contrast, place-based queries exhibit relatively strong performance across both settings, indicating that geographic knowledge is well-represented in pretraining. However, calibration improvements vary (Table 2): the person category sees highest relative ECE drop (69%), while place category shows the lowest (50%). This suggests that structured choices help correct overconfidence in ambiguous queries but offer limited calibration gains when models already retrieve knowledge with high confidence. These results demonstrate that miscalibration depends on both task framing and knowledge representation, not just model scale. While structured reasoning improves confidence alignment in person-based queries, factual retrieval tasks like place-based questions may require alternative calibration strategies to prevent persistent overconfidence.
Table 2: ECE comparison across question types.
@@ 3 Results Analysis and Discussion
4o-mini 0.77 0.32 0.73 0.35 0.76 0.25 0.73 0.42
4-turbo 0.62 0.23 0.65 0.26 0.62 0.03 0.52 0.19
4o 0.41 0.05 0.53 0.13 0.51 0.07 0.35 0.04
LLaMA3.1 0.84 0.37 0.82 0.38 0.80 0.31 0.73 0.40
LLaMA3 0.85 0.35 0.82 0.36 0.81 0.33 0.73 0.45
Gemma2 0.80 0.33 0.78 0.42 0.84 0.33 0.76 0.41
mean 0.72 0.28 0.72 0.32 0.72 0.22 0.64 0.32
Implications for Model Deployment
Our findings indicate that calibration errors are not uniform and vary based on both model size and input structure. The susceptibility of bigger models to distractors suggests that prompt engineering strategies should include explicit calibration constraints—for example, by encouraging models to justify their confidence levels rather than relying on raw probability scores. For smaller models, answer ranking or re-scoring techniques may provide more robust confidence adjustments, ensuring that they do not simply default to high-confidence selections when uncertain. Future research should explore task-specific calibration techniques rather than applying one-size-fits-all solutions, particularly for applications where overconfidence poses significant risks (e.g., legal and medical AI systems).
@@ 4 Conclusion
This study provides an investigation of calibration in LLMs, revealing key trade-offs between model size, answer format, and confidence alignment. While bigger models exhibit better calibration, they remain susceptible to distraction, whereas smaller models show dramatic performance improvements with structured answer choices but fail to meaningfully estimate uncertainty. Unlike prior work that broadly characterizes LLM miscalibration, our study pinpoints conditions under which calibration errors arise, offering a framework for mitigating overconfidence in real-world applications. Future work should explore calibration-aware fine-tuning strategies and uncertainty-aware response generation to enhance LLM reliability, particularly in high-stakes domains.
5 Limitations
Our study is limited by its focus on a small set of LLMs, excluding domain-specific or fine-tuned variants, which may affect generalizability. Additionally, the use of LLMs to pick from multiple-choice options introduces potential selection bias, as highlighted by recent work Zheng et al. (2023). While we mitigate this by randomizing option order, it may not fully eliminate bias, impacting generation reliability.
Furthermore, our reliance on GPT-4o-mini as the primary evaluation judge could be questioned, as automated LLM-based scoring may introduce biases. Although we conducted a small-scale human evaluation to validate its reliability, finding substantial agreement with human annotators (Cohen’s kappa = 0.82), minor inconsistencies were observed in ambiguous cases (see Appendix B for details).
@@ 85. The answers should be realistic and not obviously wrong
Table 3: Performance metrics of LLMs in the Normal ($\mathcal{N}$) and Distractor ($\mathcal{D}$) settings, including accuracy (correct), non-attempt (na), ECE, and the number of helped ($\mathcal{D}_{helped}$) and harmed ($\mathcal{D}_{harmed}$) instances.
@@ Appendix B Same LLM judge as Prediction LLM
We employ the same LLM as both the judge and the base model responsible for predicting answers to the questions. The performance metrics are presented in Table 3. Upon manually inspecting instances from smaller LLMs, we observe that the LLM judge occasionally misclassifies responses and refrains from assigning NOT_ATTEMPTED to certain data points. This can be seen by comparing $\mathcal{N}_{none}$ in Table1 and 3. To address this issue and ensure consistency, we use gpt-4o-mini as the LLM judge across all models. We also create reliability diagrams of pipeline where LLM-judge was different and the graphs are shown in Figure 4.
@@
Figure 4: Reliability diagrams (RDs) showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings. The numbers on top of bars represent the number of correctly predicted instances. (y-axis: actual accuracy, x-axis: predicted confidence)
Appendix C LLMs details
In our study, we utilize six distinct LLMs representing a diverse range of capabilities and architectures. The GPT models (GPT-4o-mini, GPT-4-turbo, and GPT-4o) are state-of-the-art LLMs provided via OpenAI’s API. These models are known for their advanced reasoning, contextual understanding, and robust performance across a variety of tasks. The larger models, such as GPT-4o, excel in handling complex queries and demonstrate superior accuracy, while the smaller variant, GPT-4o-mini, serves as a lightweight alternative with reduced computational overhead. We use GroqCloud for the LLaMA3.1-8B-instant and LLaMA3-8B-8192 models. These models are smaller in scale but are effective in tasks requiring efficient knowledge retrieval and basic reasoning. Finally, Gemma2-9B-it, also from GroqCloud, offers a balance between size and capability, though its performance is limited by its relatively smaller model size and less extensive pretraining. Together, these models provide a comprehensive landscape for evaluating calibration, accuracy, and decision-making across diverse architectures and computational setups.

### v2

@@ 3.5 Evaluation Metrics:
To evaluate performance, we measure correctly answered questions for both variations ($\mathcal{N}$ and $\mathcal{D}$). For calibration assessment, we use ECE to quantify the misalignment between a model’s predicted confidence and actual accuracy. A well-calibrated model produces confidence estimates that closely match its true correctness, with an ECE of zero indicating perfect calibration. Following Naeini et al. (2015), we compute ECE using empirical binning (bin size 0.1) to ensure a robust measurement of miscalibration. Additionally, we define two complementary metrics: $\mathcal{D}_{helped}$, denoting instances where the model failed under the $\mathcal{N}$ setting but succeeded when distractors were added ($\mathcal{N}_{i}=0$ and $\mathcal{D}_{i}=1$); and $\mathcal{D}_{harmed}$, capturing the reverse—cases where the model initially answered correctly under $\mathcal{N}$ but erred when distractors were introduced ($\mathcal{N}_{i}=1$ and $\mathcal{D}_{i}=0$).
4 Experimental Results and Analysis
4.1 Quantifying Baseline Calibration of LLMs
We first quantify each model’s out-of-the-box accuracy ($\mathcal{N}_{correct}$), NOT_ATTEMPTED ($\mathcal{N}_{na}$), and $\mathcal{N}_{ECE}$ on three benchmarks: SimpleQA (a deliberately hard, concise factoid task), FaVIQ (moderate difficulty), and TriviaQA (an easier, QA dataset). Table 1 reports the full metrics.
On the hardest SimpleQA benchmark—where direct “off-the-shelf” generation is inherently complex—even GPT-4o attains only 35% accuracy (ECE 0.45, NOT_ATTEMPTED $\approx$ 8%). Larger GPT-4 variants have been pretrained on vastly more tokens (and may even have encountered similar questions), yet their calibration loss remains on the same order as smaller models. This parity—despite scale and pretraining volume—indicates persistent overconfidence across sizes on challenging questions.
By contrast, the easy TriviaQA setting reveals the limits of overconfidence: GPT-4o’s accuracy rises to 90% with ECE $\approx$ 0.07 and near-zero NOT_ATTEMPTED, while smaller or open-source models tighten their reliability curves more markedly. In other words, on simpler, context-rich queries, larger models not only answer correctly but also exhibit proportionally less overconfidence compared to hard benchmarks. FaVIQ again falls between these extremes, with both accuracy and ECE interpolating smoothly.
Comparative analysis across families shows that GPT-4 variants consistently occupy the “high accuracy, low ECE, low NOT_ATTEMPTED” regime on all datasets. Small LLaMA-3 models, in contrast, post single-digit accuracy on SimpleQA, ECEs approaching 0.8, and frequent deferrals; scaling them to 70B or adopting to other open-source alternatives (Gemma2-9b-it, Qwen-qwq-32b) yields only incremental gains.
4.2 Effects of Structured Distractors on Accuracy and Confidence
To quantify effect of adding correct answer alongside three incorrect options, we measure relative accuracy gain $\Delta\mathrm{Acc}=(\mathcal{D}_{\mathrm{correct}}-\mathcal{N}_{\mathrm{correct}})/\mathcal{N}_{\mathrm{correct}}$ and ECE compression $\Delta\mathrm{ECE}=\mathcal{N}_{\mathrm{ECE}}-\mathcal{D}_{\mathrm{ECE}}$. Table 1 reports counts and percentage of $\mathcal{D}_{\mathrm{helped}}$ vs. $\mathcal{D}_{\mathrm{harmed}}$ and NOT_ATTEMPTED ($\mathcal{D}_{\text{na}}$). Figure 2 overlays reliability diagrams (RDs) for all the nine models on SimpleQA dataset. RDs for FaVIQ and TriviaQA are in the Appendix E.
[…]
Smaller models (< 10B) show the largest $\Delta\mathrm{Acc}$ on FaVIQ and SimpleQA but also the highest $\mathcal{D}_{\mathrm{harmed}}$ rates on FaVIQ and TriviaQA, suggesting that limited pretraining makes them more susceptible to distractor-induced errors. Notably, $\mathcal{D}_{\text{na}}$ drops for all models (to zero on TriviaQA), underlining that no LLM abstains once explicit options are provided for easier questions.
4.3 Influence of Fine-Tuning Regime and Model Architecture
Our analysis systematically investigates how different fine-tuning strategies and model architectures impact accuracy and calibration performance, specifically in the context of structured distractors.
First, we observe that effectiveness of RLHF on calibration performance varies significantly across different model implementations and sizes. While RLHF models such as GPT-4o-mini (assumed  8B parameters) exhibit superior calibration performance (ECE 0.750 reduced to 0.320 on SimpleQA), smaller-scale RLHF models like LLaMA-3-8b-8192 and LLaMA-3.1-8b-Instant underperform relative to Qwen-qwq-32B, an SFT-only model. Specifically, in $\mathcal{D}$ setting, Qwen-qwq-32B demonstrates better accuracy (51.68% vs. 44.64% for LLaMA-3.1 and 44.01% for LLaMA-3-8b on SimpleQA) and calibration (ECE 0.253 vs. 0.367 for LLaMA-3.1 and 0.361 for LLaMA3-8b), highlighting that RLHF alone does not guarantee superior calibration. This indicates that other factors, such as the volume and diversity of training data, the quality of fine-tuning data, and overall training strategies, play crucial roles.
Figure 3: Accuracy and calibration shifts with distractors. We show relative accuracy gains (bars) and ECE changes (points) when distractor options are added. While all models improve in accuracy, calibration effects vary—large models benefit most, while smaller or models often remain miscalibrated.
Examining within the LLaMA family, LLaMA 3.1–8b–Instruct notably outperforms the earlier LLaMA3-8B variant across all benchmarks despite identical parameter counts. This improved performance, especially evident in all three open-domain QA tasks in $\mathcal{N}$ setting (e.g. accuracy 75.46% vs. 67.41% in TriviaQA dataset), stems from enhancements in parametric knowledge, refined instruction tuning, and superior calibration. LLaMA-3.1-8b benefited from training on an extended and more recent dataset (up to December 2023), enabling better retention of long-tail factual knowledge and improved instruction-following capabilities, thereby enhancing robustness and reliability.
Distilled models, notably Gemma2-9b-it, exhibit higher "NOT_ATTEMPTED" rates on all the three datasets (e.g. 33.29% on SimpleQA) compared to similarly sized models, indicating challenges in effectively utilizing their compressed knowledge base without external support. Qwen-qwq-32b, despite lacking RLHF fine-tuning, consistently produces answers with lower "NOT_ATTEMPTED" rates and demonstrates robustness against distractor-induced errors, as indicated by its lower percentage of harmed instances.
Furthermore, although significantly larger, the MoE-based LLaMA-4-Scout-17b does not outperform GPT-4o-mini in accuracy or calibration, underscoring that training volume and quality significantly impact performance more than sheer parameter count. However, at a substantially larger scale, LLaMA-3-70b successfully surpasses GPT-4o-mini, highlighting the interplay of extensive parameterization, robust training, and fine-tuning strategies.
In conclusion, our comprehensive analysis reveals that while RLHF can substantially improve model calibration, it is not universally effective without careful consideration of other critical training factors. Models leveraging large-scale training datasets, updated fine-tuning approaches, and comprehensive instruction tuning achieve optimal accuracy and calibration. Effective deployment in reliability-sensitive contexts thus demands a strategic blend of parameterization, extensive pre-training, robust fine-tuning methodologies, and carefully designed calibration interventions.
@@ 4.3 Influence of Fine-Tuning Regime and Model Architecture
GPT-4o-mini 0.77 0.32 0.73 0.35 0.76 0.25 0.73 0.42
GPT-4-turbo 0.62 0.23 0.65 0.26 0.62 0.03 0.52 0.19
GPT-4o 0.41 0.05 0.53 0.13 0.51 0.07 0.35 0.04
LLaMA-3.1-8b 0.84 0.37 0.82 0.38 0.80 0.31 0.73 0.40
LLaMA-3-8b 0.85 0.35 0.82 0.36 0.81 0.33 0.73 0.45
LLaMA-3-70b 0.80 0.22 0.80 0.28 0.76 0.19 0.65 0.33
LLaMA-4-17b 0.68 0.30 0.65 0.29 0.60 0.24 0.53 0.35
Gemma2-9b-it 0.80 0.33 0.78 0.42 0.84 0.33 0.76 0.41
Qwen-qwq-32b 0.68 0.21 0.66 0.25 0.68 0.26 0.66 0.31
mean 0.72 0.26 0.72 0.30 0.71 0.22 0.63 0.32
(a) ECE comparison across question types
(b) Date
(c) Number
(d) Person
(e) Place
4.4 Effect of Model Size within LLM Families
We disentangle parameter count from other factors by comparing the smallest, mid-sized, and largest checkpoints released by each provider. In the normal setting ($\mathcal{N}$), accuracy increases monotonically with scale for both families; however, calibration improves much faster for the OpenAI series. GPT-4o already attains an ECE of 0.450 on SimpleQA. This suggests that RLHF alignment, used uniformly across GPT-4 models, amplifies the natural size-driven gains in self-assessment that emerge from scaling alone.
Introducing distractor options ($\mathcal{D}$) radically alters the picture. The largest relative accuracy jumps occur in the smallest models as shown in Figure 3: GPT-4o-mini rockets from 8.5% to 47.4% accuracy on SimpleQA (+461%), and LLaMA-3-8b leaps +819% over the same split. In contrast, their flagship counterparts—GPT-4o and LLaMA-70b—gain a more modest +109% and +338%, respectively. Yet these headline boosts do not translate into equally dramatic calibration improvements. After distractors, GPT-4o compresses its ECE by 92% (to 0.037), whereas GPT-4o-mini still lingers above 0.32. A mirror pattern holds for LLaMA: the 70B model cuts ECE by 69%, finishing at 0.24, while the 8B base remains mis-calibrated (ECE 0.36) despite its vast accuracy lift. When we extend this analysis to FaVIQ and TriviaQA, the same ranking by model size holds but with attenuated returns. On FaVIQ, distractors yield solid—but more moderate—accuracy boosts across all scales. On TriviaQA, where the baseline performance is already high, relative gains shrink into the low-teens, indicating only marginal benefit from explicit distractors.
Two key takeaways emerge from our analysis. First, small models acquire factual knowledge more quickly than they learn to assess their own confidence: adding explicit answer choices boosts accuracy but yields poorly calibrated probability estimates. Second, increasing model scale primarily improves a model’s ability to quantify its certainty rather than uncover new knowledge. Consequently, when downstream tasks rely directly on confidence scores—such as risk-aware planning or answer adjudication—larger models offer more trustworthy probabilities. In contrast, in settings where latency or compute cost is the main constraint and some post-hoc temperature scaling is acceptable, smaller models can still deliver adequate confidence estimates so long as they’re provided with distractors or retrieval-augmented context.
4.5 Performance Across Question Types
The SimpleQA dataset already categorizes 4326 questions into four non-overlapping types—Date (1418), Number (663), Person (1041), and Place (427)—and we evaluate each model’s accuracy and ECE in both the free-generation ($\mathcal{N}$) and distractor-augmented ($\mathcal{D}$) settings. To better understand calibration weaknesses, we analyze model performance and confidence alignment across these question types (Figure 4). Person-based queries are most challenging, likely due to name ambiguities and inherent variability in names, overlapping roles, and contextual dependencies that require deeper reasoning beyond surface-level pattern matching. LLMs frequently confuse historical figures with similar names, but providing structured answer choices significantly improves accuracy, suggesting that explicit disambiguation helps mitigate uncertainty in this category. In contrast, place-based queries exhibit relatively strong performance across both settings, indicating that geographic knowledge is well-represented in pretraining. However, calibration improvements vary (Figure 4(a)): the person category sees highest relative ECE drop (69%), while place category shows the lowest (49%). This suggests that structured choices help correct overconfidence in ambiguous queries but offer limited calibration gains when models already retrieve knowledge with high confidence. These results demonstrate that miscalibration depends on both task framing and knowledge representation, not just model scale or architecture. While structured reasoning improves confidence alignment in person-based queries, factual retrieval tasks like place-based questions may require alternative calibration strategies to prevent persistent overconfidence.
@@ 5 Conclusion
Our investigation provides a rigorous empirical foundation for understanding and addressing calibration issues in LLMs. We reveal widespread overconfidence across various model families and sizes, significantly improved by employing structured distractors—particularly effective for smaller models. However, our findings also highlight counterintuitive outcomes, including degraded calibration in large models on simpler queries. Moreover, systematic miscalibration across specific query categories underscores the complexity of the calibration challenge beyond simple accuracy. Consequently, achieving trustworthy AI requires a multifaceted calibration strategy, integrating robust RLHF, optimized prompt design, and post-hoc calibration adjustments. The evaluation framework and guidelines proposed herein serve as critical tools for future research, driving forward the development of LLMs that are not only accurate but reliably calibrated for safe real-world application.
Limitations
While our work breaks new ground in structured distractor–augmented calibration, we note three main limitations:
Model & task scope.
We cover nine state-of-the-art LLMs across three single-turn factoid QA benchmarks, which gives a broad but still focused view of calibration behavior. Extending to domain-specialized models, multi-hop reasoning, and generative tasks is an important direction for future validation.
Distractor realism.
To ensure consistency, all distractors were produced by a controlled LLM prompt. Although this may not capture every nuance of real-world retrieval, we followed strict prompting guidelines and shuffled options to minimize artefacts Zheng et al. (2023)—nevertheless, evaluating with human-curated or retrieval-sourced distractors remains valuable future work.
Automated-judge bias.
We use GPT-4o-mini as our primary correctness judge, complemented by a human evaluation (Cohen’s $\kappa$ = 0.82, see Appendix B for details) to confirm its reliability. Yet, subtle misclassifications could still subtly affect accuracy metrics, so applying complementary human or ensemble-based judges would further strengthen the results.
@@ 85. The answers should be realistic and not obviously wrong
Table 2: Performance metrics of LLMs on SimpleQA dataset in the Normal ($\mathcal{N}$) and Distractor ($\mathcal{D}$) settings, including accuracy (correct), non-attempt (na), ECE, and the number of helped ($\mathcal{D}_{helped}$) and harmed ($\mathcal{D}_{harmed}$) instances. Here the LLM judge model is same as the prediction model.
@@ Appendix B Same LLM judge as Prediction LLM
We employ the same LLM as both the judge and the base model responsible for predicting answers to the questions. The performance metrics are presented in Table 2. Upon manually inspecting instances from smaller LLMs, we observe that the LLM judge occasionally misclassifies responses and refrains from assigning NOT_ATTEMPTED to certain data points. This can be seen by comparing $\mathcal{N}_{none}$ in Table1 and 2 for the SimpleQA dataset. To address this issue and ensure consistency, we use GPT-4o-mini as the LLM judge across all models. We also create reliability diagrams of pipeline where LLM-judge was different and the graphs are shown in Figure 5.
@@
Figure 5: Reliability diagrams (RDs) on SimpleQA dataset showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings. The numbers on top of bars represent the number of correctly predicted instances (y-axis: actual accuracy, x-axis: predicted confidence). Here the LLM judge model is same as the prediction model.
Appendix C Distractor Generation
Our distractors were generated using GPT-4o-mini with a carefully designed prompt ensuring distractors were factually incorrect yet plausible (matching answer type, specificity, and context). To validate distractor plausibility further, we conducted a manual inspection on a subset (500+ randomly examples) from SimpleQA dataset, confirming that distractors were consistently plausible & contextually relevant.
Appendix D LLMs details
Table D surveys nine leading LLMs, revealing how differences in scale, architecture, and pretraining data shape their alignment pipelines. The GPT-4 family and LLaMA-3 series are dense models trained on massive multimodal or text-only corpora, each undergoing both SFT and RLHF before instruct-tuning. In contrast, LLaMA-4-Scout employs a 16-expert MoE design over mixed vision-text data, while Gemma2-9b-it and Qwen-qwq-32b explore distillation-based and pure SFT regimes, respectively. Despite these varied strategies, all nine models receive dedicated fine-tuning and instruction-tuning to optimize performance and calibration in downstream QA and conversational tasks.
Table 3: Taxonomy of selected LLMs showing differences in training and fine-tuning approaches, where FT (fine-tuning) and IT (instruct-tuning).
Model (Params / Context) Architecture Dataset Type Training Strategy FT IT
GPT-4o
(undisclosed / 128K)
Dense Multi-modal (web text, code, images, audio transcripts) Pre-training + RLHF
GPT-4o-mini
(undisclosed / 128K)
Dense Multi-modal (web text, code, images) SFT + RLHF
GPT-4-turbo
(undisclosed / 128K)
Dense Multi-modal (web text, code, images) Pre-training + SFT + RLHF
LLaMA-3.1-8B-Instant
(8B / 128K)
Dense (GQA) 15T tokens – Public (web, code, multilingual) Pre-training + SFT + RLHF
LLaMA-3-8B-Instruct
(8B / 8K)
Dense (GQA) 15T tokens – Public (web, code, multilingual) Pre-training + SFT + RLHF
LLaMA-3-70B-Instruct
(70B / 8K)
Dense (GQA) 15T tokens – Public (web, code, multilingual) Pre-training + SFT + RLHF
LLaMA-4-Scout-17B
(17B / 10M)
MoE (16 experts) 40T tokens – Mixed (text + vision, multilingual) MoE Pre-training + SFT + RLHF
Gemma2-9B-it
(9B / 8K)
Dense (GQA, interleaved local–global) 8T tokens – Public (web, academic) Distillation + SFT + RLHF
Qwen-qwq-32B
(32B / 131K)
Dense (GQA) 18T tokens – Public (multilingual; web text, code, scientific lit.) Pre-training + SFT
Appendix E Reliability Diagrams of Selected Models on FaVIQ and TriviaQA dataset
Figure 6: Reliability diagrams (RDs) showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings on the TriviaQA dataset. (y-axis: actual accuracy, x-axis: predicted confidence).
Figure 7: Reliability diagrams (RDs) showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings on the FaVIQ dataset. (y-axis: actual accuracy, x-axis: predicted confidence)